OpenStudy (ikou):

Please help me Simplify square root 24 minus 3 times the square root of 45 plus 2 times the square root of 20.

OpenStudy (ikou):

https://gyazo.com/0337b4f702f3e8ad722458f8e4f85956 equation

OpenStudy (ikou):

choices https://gyazo.com/2a4bc47d0c0d6ba17ba947765c6668b4

OpenStudy (michele_laino):

hint: we have to rewrite the expression, using similar radicals

OpenStudy (michele_laino):

we can write this: \[\Large \begin{gathered} \sqrt {24} = \sqrt {6 \cdot {2^2}} = 2\sqrt 6 \hfill \\ \sqrt {45} = \sqrt {{3^2} \cdot 5} = 3\sqrt 5 \hfill \\ \sqrt {20} = \sqrt {{2^2} \cdot 5} = 2\sqrt 5 \hfill \\ \end{gathered} \]

OpenStudy (ikou):

So would it be c?

OpenStudy (michele_laino):

no, I'm sorry Please, if I substitute, the above simplification, into your expression, I can write: \[\huge \begin{gathered} \sqrt {24} - 3\sqrt {45} + 2\sqrt {20} = \hfill \\ \hfill \\ = 2\sqrt 6 - 3\left( {3\sqrt 5 } \right) + 2\left( {2\sqrt 5 } \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (ikou):

hmm.. this is hard for me sorry. I see B

OpenStudy (michele_laino):

that's right!

OpenStudy (ikou):

oh awesome, thank you for helping me out!

OpenStudy (michele_laino):

:)